13. Calculate equilibrium concentrations: [CO2] 0.56 CO2(8)+ C(graphite) 2 CO(g) M K=2.25
Answer #1

                     CO2 + C (Graphite) ----------> 2 CO

Initially         0.56        0                               0

At equili      (0.56 - x)   x                              2x

Keq = [CO]^2/[CO2][C]

2.25 = (2x)^2*/((0.56 - x)(x))

x = 0.2016

At equilibrium,

[CO2] = 0.56 - 0.2016 = 0.3584 M

C(graphite) = 0.2016 M

[CO] = 2x = 2*0.2016 = 0.4032 M